OpenStudy (anonymous):

A circular test track for cars has a circumference of 2.7km . A car travels around the track from the southernmost point to the northernmost point. What distance does the car travel? (answer 1.35km). Part B !!!!!!!What is the car's displacement from its original position?

OpenStudy (anonymous):

It means it traveled distance equal to semicircle of the circle...

OpenStudy (anonymous):

displacement will be along the diameter of the circle.

OpenStudy (anonymous):

see the circumference of circle is given by \[2\pi r\] and thus distance travelled is semicircle so \[2\pi r=2.7\] \[\pi r=2.7/2=1.35km\]

OpenStudy (anonymous):

Yah I got the first part. Which was 1.35km but the displacement is asking for the same thing?

OpenStudy (anonymous):

Not sure what you are trying to explain.

OpenStudy (anonymous):

ok so as i said displacement is shortest path travelled. In case of circle what os shoretest path traveled from soth point to the northpoint? It's diameter right. @f_s32

OpenStudy (anonymous):

just you need to calculate diameter of the circle :)